OpenStudy (anonymous):

Dante is making a necklace with 18 rows of tiny beads in which the number of beads per row is given by the series 3+10+17+24+.... If you were to write this series in summation notation, give... 1. The lower limit of the sum 2. The upper limit of the sum and 3. The explicit formula of the sum.

OpenStudy (fortytherapper):

The lower limit is in the form n = ?, where the ? is the beginning number. Do you see the beginning number? It's given actually The upper limit is just a number, but it's the ending number. They gave this one too, if you see it?

OpenStudy (anonymous):

Oh okay, so the lower limit is 3 and the upper limit is 24??

OpenStudy (anonymous):

Is that correct?

OpenStudy (fortytherapper):

I didn't even know OpenStudy had this! \[\sum_{n = ?}^{ending}Formula\] But the answer lie within the row of beads. How many beads do we start with and how many do we end with?

OpenStudy (anonymous):

So 3 to infinity? I'm sorry, this question just confuses me.

OpenStudy (fortytherapper):

No worries Say you wanted to make 18 rows of beads like Dante. How many rows (besides 0) would you start with and how many rows would you end with?

OpenStudy (welshfella):

this is an arithmetic sequence with common difference 7. 7 is added to get the next term

OpenStudy (anonymous):

oh lol start with 1 end with 18

OpenStudy (anonymous):

Yeah welsh, I see that...

OpenStudy (fortytherapper):

Right, so let's add those in \[\sum_{n=1}^{18}Formula\]

OpenStudy (welshfella):

nthe term = a1 + (n -1)d where a1 = first term and d = common difference so the last (18th term) is 3 + (18-1)7

OpenStudy (fortytherapper):

Since Row 1 = 3 Row 2 = 10 Row 3 = 17 etc., now we wanna make a formula for it

OpenStudy (welshfella):

sum of n terms = (n/2)[a1 + L] where a1 = first term , L = last term) n = 18 , a1 = 3 and L is 3 + 17*7

OpenStudy (anonymous):

The pattern ends in 122

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

So would that be the upper limit?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

So wait, how do I find the formula?

OpenStudy (welshfella):

Oh they want summation notation . ......

OpenStudy (anonymous):

yes

OpenStudy (welshfella):

excuse me!!

OpenStudy (anonymous):

???

OpenStudy (welshfella):



OpenStudy (welshfella):

the limits are 1 and 18

OpenStudy (anonymous):

Just f(n)? So would it be 1(7)??

OpenStudy (welshfella):

and the expicit fformula for each term is a1 + (n - 1)d

OpenStudy (anonymous):

So 3+(18-1)7

OpenStudy (welshfella):